# Can AI be enabled to dynamical downscaling? Training a Latent Diffusion Model to mimic km-scale COSMO-CLM downscaling of ERA5 over Italy

### v1

@@ 4.1 UNET Architecture
The core UNET network implemented for our experiments is a standard UNET architecture ([40]), featuring an encoder (contracting path), a bottleneck, and a decoder (expansive path), with skip connections bridging corresponding levels between the encoder and decoder to preserve spatial information. To use a standard UNET to perform downscaling, the input low-resolution data are interpolated with the nearest neighbor interpolation to the target high resolution before feeding them to the network. The encoder is composed of four blocks, each consisting of a layer containing two consecutive 2D convolutions with ReLU activation, interspersed with batch normalization to ensure stable learning ([41]), and a max-pooling operation. The max-pooling layer reduces the spatial resolution by half, enabling the model to capture increasingly complex features while reducing the image dimensions. The output of each encoder block is used in both the next encoder block and the corresponding decoder block through skip connections. The decoder mirrors the encoder with transposed 2D convolutional layers and upsampling steps to go back to the starting resolution. The use of batch normalization ensures robust learning, while the skip connections help preserve critical spatial information across the encoder-decoder bridge. Details on the UNET structure and resolutions are depicted in Figure 3.
@@ 4.2 GAN Architecture
The Generative Adversarial Network (GAN, [19, 20]) tested in this experiment consists of deep, fully convolutional Generator and Discriminator networks, conditioned with low-resolution predictors and high-resolution static data. The generator is trained to output fields that cannot be distinguished from ground truth images by a discriminator, which is trained on the other hand to detect the generator’s ’fake’ outputs. Our reference GAN consists of a UNET generator upgraded with a Patch-GAN discriminator ([42]). The input data to the generator are low-resolution predictors and high-resolution static data only (no noise addition is performed) and we, therefore, obtain a deterministic GAN.
@@
The discriminator is a PatchGAN convolutional classifier ([42]), which focuses on structures at the scale of image patches. The structure of the discriminator is composed of modules of the form convolution-BatchNorm-ReLu. It assigns a "realness" score to each N × N patch of the image, runs convolutionally across the image, and allows us to obtain an overall score by averaging all responses for each patch. High-quality results can be obtained with patches much smaller than the full size of the image, with relevant advantages in terms of resources for the training and application to arbitrarily large images. Details on the discriminator network structure and resolutions are depicted in Figure 3.
The training procedure follows the combined loss function approach for GANs [20], including recent improvements to promote stability in the training ([43]), with the primary goal of balancing the minimization of both the generator’s and discriminator’s losses, which are adversarial. The pixel loss we used is the mean absolute error, while the discriminator loss is the hinge loss. The discriminator is activated after 50000 training steps, giving the generator time to learn to generate consistent outputs and thus stabilize the adversarial training ([43]). After activating the discriminator, the network is trained by updating alternatively the gradients of the generator and the discriminator. The network is constantly fed with the whole target domain and no patch-training is applied.
@@ 4.3 Latent Diffusion Model Architecture
Diffusion Models ([21]) are probabilistic models meant to extrapolate a data distribution p(x) by corrupting the training data through the successive addition of Gaussian noise (fixed) and then learning to recover the data by reversing this noising process (generative).
The Latent Diffusion Model (LDM) applied for this experiment is an architecture derived from Stable Diffusion ([24]), specifically a re-adaptation of the conditional LDM LDCast ([30]), developed for precipitation nowcasting and already successfully applied for other variables (e.g. [44]). The latent diffusion model derived for the downscaling task is composed of three main elements: (i) a convolutional Variational AutoEncoder (VAE), used to project the residual high-resolution target variables to and from a latent space; (ii) a conditioner, used to embed low-resolution data and high-resolution static data; and a denoiser, used to manage the diffusion process within the latent space. In the following sections, we report a detailed description of each model component. The number of trainable parameters for each component is summarized in Table  1, while Figures 4 and 5 summarize the training and inference procedures for the model and the main structure of each component architecture, respectively.
@@ 4.3.1 Variational Autoencoder
The encoder and the decoder are structured as 2D convolutional networks composed of blocks of a ResNet residual block ([45]) and a downsampling/upsampling convolutional layer. Three levels of such blocks are used, each reducing each spatial dimension by a factor of 2, while the number of channels is bottlenecked at 32 times the number of input target variables (i.e. 32x1 for the 2-m temperature and 32x2 for the 10-m wind speed components). The VAE bottleneck latent space is regularised with a loss based on Kullback-Leibler divergence (KL) between the latent variable and a multivariate standard normal variable.
While the space dimensions are reduced by a factor of $8^{2}$ (from 512x512 to 64x64 pixel patches), the number of channels is increased from 1 (for the 2-m temperature) and 2 (for the 10-m wind speed components) to 32 and 64, respectively: the overall amount of data is therefore compressed only by a factor of 2, for both the VAEs (from 1x512x512 to 32x64x64 and from 2x512x512 to 64x64x64). Nevertheless, the gain in training performance of the denoiser and conditioner is much greater than the data reduction factor as the compression along the space dimension is more relevant for reducing the computational cost of the training than the increase in channel number ([24]).
@@ 4.3.2 Conditioner
The conditioner stack acts as a context encoder to process the low-resolution predictors and high-resolution static data and embed them into each level of the denoiser UNET architecture. Initially, both datasets are preprocessed by passing through a dedicated encoder, a projection layer, and an analysis sequence, before being merged. The predictors’ encoder is a basic Identity layer as they already have the same spatial dimensions as the latent space (64x64). The static data encoder is a variational encoder with the same structure as the VAE described in the previous Section 4.3.1. For both datasets, the projection layer is a 2D convolutional layer with unitary kernel size, used to increase the number of channels, and the analysis is a sequence of 4 2D Adaptive Fourier Neural Operator (AFNO) blocks (following [46]), used to extract relevant features. After pre-processing, the conditioning information is prepared to be fed into each level of the denoiser UNET by applying a combination of average pooling and 2D Resnet layers. Figure 4 depicts details of the conditioner’s structure.
@@ 4.3.3 Denoiser
Our denoising stack is structured as the one of the LDCast model ([30]), a re-adaptation of the U-Net-type network applied in the original latent diffusion model ([24]). The resulting denoiser network consists of a UNet backbone enabled with a conditioning mechanism based on 2D AFNO blocks ([30]), aiming at a cross-attention-like operation (as suggested in [47]). This structure is meant to control the high-resolution synthesis process feeding the conditioning in each level of the UNET architecture.
@@
To improve the reconstruction of extreme values (for both temperature and wind speed) we implemented the v-prediction parameterization in our LDM model, following [48]: this parameterization trains the denoiser to model a weighted combination of both the noise and the start image, instead of either the only noise or the only start image as done in the more traditional implementations eps and x0, respectively.
As shown in Figure 5, the conditioner and the denoiser are trained together, minimizing the mean square error (L2), and feeding the network with random patches of ERA5 predictors (64x64 pixels) and static data (512x512 pixels) for the conditioning, and high-resolution target variables (512x512 pixels) for the ground truth.
@@ 5 Results
The following sections compare the results from LDM_res against the baselines using various verification metrics and distributions. In the additional material, Section 8.4, we report the comparison of results from the LDM trained with and without the residual approach to provide an overview of the improvements introduced by this method.
@@ 5.1 Qualitative evaluation
To provide a qualitative and perceptual overview of the obtained results, we present a random snapshot of downscaled variables compared with both the input ERA5 low-resolution data and the COSMO-CLM high-resolution reference-truth (Figure 6). The second and third columns show a zoom-in on Sardinia Island, providing a deeper overview of models’ performance over complex terrain, coastal shores, and open sea. Both generative models, the GAN and LDM_res, effectively overcome the blurriness observed in both the quadratic interpolation and the UNET for the target variables. Particularly for 2-meter temperatures, LDM_res demonstrates a remarkable ability to identify and reconstruct discontinuities in the variable field (zoomed-in view in Figure 6). Figure 6 also includes results for 10-meter wind speed (in color), which is a derived field obtained by combining the two actual target variables of the models, U and V. Perceptually, the results for this variable from both the GAN and LDM_res appear similar and equally plausible, displaying significantly more small-scale features compared to the UNET. A deeper qualitative examination reveals that the GAN aligns well with the reference truth, particularly over land, but exhibits mode collapse over the sea for both target variables. An example of this effect is shown in the additional material, Section 8.1, Figure 11. Conversely, the LDM_res consistently generates plausible high-resolution data across the entire domain, over both land and sea, and for both target variables.
Figure 6: Downscaled variables from all the tested models against low-resolution ERA5 input data and high-resolution COSMO-CLM reference truth, for a randomly picked timestamp. The left columns refer to 2-m temperature, and the right columns refer to 10-m wind speed. The second and fourth columns show a zoom-in on Sardinia Island.
@@ 5.2 Verification deterministic metrics
Figure 7 compares model results for different deterministic metrics, averaging results over the whole domain for each test timestep. In addition to results from the baseline and tested models, Figure 7 also reports results for the VAE of the LDM. These results are obtained using the VAE offline, feeding it with COSMO_CLM high-resolution data and calculating the metrics on the reconstructed data: this allows quantifying the LDM error resulting from the data decompression from the latent space only. We present three distance metrics, the Root Mean Square Error (RMSE), the mean bias (BIAS), and the coefficient of determination (R2), and one correlation metric, the Pearson Correlation Coefficient (PCC). Details on how these metrics are calculated are presented as additional material in Section 8.2, following [49].
@@
The results show that the VAE contribution to LDM_res is the lowest for the RMSE of 2-m temperature, while bias, R2, and PCC have very little to no effects on temperature and wind speed.
@@ 5.3 Spatial distribution of errors
The spatial distribution of averaged-in-time magnitude differences for both the target variables and all tested models is illustrated in Figure 8. Within each panel, the numbers in squared brackets represent the 0.5 and 99.5 percentile values, offering insight into the highest errors recorded. Negative and positive values signify underestimation and overestimation, respectively, for both variables. Results from the quadratic interpolation of ERA5 data provide information on the original input data: 2-m temperature tends to be highly overestimated over complex terrain but underestimated on flat terrain, with smaller errors over the sea. Wind speed, conversely, is largely underestimated over land, particularly over mountain ridges, with a tendency towards overestimation along coastal shores.
@@
As for the wind speed results, all models tend to underestimate the wind speed. The LDM model shows the best performance, with errors reduced to nearly 0 over most of the domain, with a uniform distribution over both land and sea. The GAN presents traces of its characteristic mode collapses especially over the sea: this evidence suggests that the mode collapses are reproduced statically by the model in the same locations through time, which is consistent with the deployed training approach, i.e. feeding the network always the entire, fixed domain.
Regarding wind speed, all models exhibit a tendency to underestimation. LDM_res demonstrates superior performance, minimizing errors to nearly zero over most of the domain, with a uniform distribution over land and sea. The GAN displays traces of mode collapses, particularly evident over the sea, indicating that these collapses persist statically in specific locations over time, consistent with the training approach of presenting the network with the fixed, entire domain.
@@ 5.4 Frequency distributions
Figure 9: Comparison of frequency distributions for results from the tested models against COSMO-CLM reference-truth. The left column refers to the 2-m temperature, and the right column refers to the 10-m wind speed. Counting of pixel-wise data is cumulated for the yearly test dataset over bins of 0.5 °C and 0.05 m/s for temperature and wind speed, respectively. Notice that y-axes are logarithmic to highlight the tails of the distributions, hence the extreme values. The top row focuses on the tails of the distributions, i.e. on extreme values, while the bottom row focuses on the most frequent values and is a zoom-in on the dashed boxes for each variable.
@@
Reconstructing the distribution of 10-m wind speed proves more challenging for all models, given the inherent chaotic nature of the $u$ and $v$ wind components compared to temperature, which is strongly influenced by terrain elevation. Nonetheless, performance outcomes mirror those of the 2-m temperature. The GAN modestly improves upon UNET results, primarily in reducing occurrences of low wind speeds. LDM_res exhibits the highest proficiency in capturing both the tail and center of the wind speed distribution.
@@ 5.5 Radially Averaged Power Spectral Density (RAPSD)
Figure 10: Comparison of Radially Averaged Power Spectral Density (RAPSD) distributions for results from the tested models against COSMO-CLM reference-truth. The left column refers to the 2-m temperature, and the right column refers to the 10-m wind speed. The first row shows the averaged-in-time spectra, across the whole test dataset. Notice that in the first row y-axes are logarithmic to highlight the tail of the distributions, hence the high frequencies. The bottom rows show the distributions of single-time RAPSD values for fixed wavelengths, namely 269, 20, 9, and 5 km.
@@
Figure 10 showcases the results in terms of Radially Averaged Power Spectral Density (RAPSD), computed following the implementation outlined in [50]. The top row of the figure illustrates a single RAPSD, representing the average of each RAPSD calculated for every timestamp within the test dataset. To provide insight into the distribution of these values across all timestamps, the distributions of single-time RAPSD for fixed wavelengths are displayed in the bottom rows of Figure 10.
@@
Overall, all DL models effectively reconstruct the 2-m temperature power spectra down to wavelengths of 10 km. However, LDM_res consistently outperforms both the UNET and the GAN, as evident in panels (b) and (c) of Figure 10. The UNET and the GAN yield similar results, with marginal yet consistent enhancements stemming from the adversarial training of the UNET. The diffusion process of LDM_res adeptly enhances the generation of small-scale features, showing precise reconstruction of the spectra up to 9 km (see panel c). For scales smaller than 9-10 km, all models exhibit decreased performance, albeit still showing improvements over the quadratic interpolation of ERA5. LDM_res still outperforms the other models but the very small-scale variability of the original data is slightly underestimated. This behavior is to be ascribed to the VAE, as further elucidated in Section 8.3. Indeed, original reference-truth data compressed and reconstructed by the VAE show the very same spectra underestimation for scales smaller than 9 km. The loss of information is therefore due to and inherent to the projection to the latent space.
@@
In contrast, results for wind speed distinctly demonstrate that generative models surpass both quadratic interpolation and UNET, effectively matching the slope of the energy power spectra and remaining competitive with each other. Specifically, LDM_res consistently outperforms the GAN up to 7 km, as emphasized in panels (b) and (c) of Figure 10 (right column). Similar to the 2-m temperature, for scales smaller than 10-9 km, both the GAN and LDM_res experience reduced performance, although they consistently exhibit improvements over the UNET. This behavior, for LDM_res, is in this case partly to be ascribed to the VAE, as further shown in Section 8.3. Original wind reference-truth data compressed and reconstructed by the VAE show a similar spectra underestimation for scales smaller than 9 km but an additional loss is to be attributed intrinsically to the extraction of features with the diffusion process conditioned with the low-resolution data and high-resolution static data. The loss of information is therefore due to and inherent to the projection to the latent space.

### v2

@@ 3.1 UNET Architecture
The core UNET network implemented for our experiments is a standard UNET architecture (Ronneberger et al., 2015), featuring an encoder (contracting path), a bottleneck, and a decoder (expansive path), with skip connections bridging corresponding levels between the encoder and decoder to preserve spatial information. To use a standard UNET to perform downscaling, the input low-resolution data are interpolated with the nearest neighbor interpolation to the target high resolution before feeding them to the network. The encoder is composed of four blocks, each consisting of a layer containing two consecutive 2D convolutions with ReLU activation, interspersed with batch normalization to ensure stable learning (Ioffe and Szegedy, 2015), and a max-pooling operation. The max-pooling layer reduces the spatial resolution by half, enabling the model to capture increasingly complex features while reducing the image dimensions. The output of each encoder block is used in both the next encoder block and the corresponding decoder block through skip connections. The decoder mirrors the encoder with transposed 2D convolutional layers and upsampling steps to go back to the starting resolution. The use of batch normalization ensures robust learning, while the skip connections help preserve critical spatial information across the encoder-decoder bridge. The total number of trainable parameters for the UNET is $\sim$31 M (Table 1). Details on the UNET structure and resolutions are depicted in Figure 3.
@@ 3.2 GAN Architecture
The Generative Adversarial Network (GAN) (Goodfellow et al., 2014, 2020) tested in this experiment consists of deep, fully convolutional Generator and Discriminator networks, conditioned with low-resolution predictors and high-resolution static data. The generator is trained to output fields that cannot be distinguished from ground truth images by a discriminator, which is trained on the other hand to detect the generator’s ’fake’ outputs. Our reference GAN consists of a UNET generator upgraded with a Patch-GAN discriminator (Isola et al., 2017). The input data to the generator are low-resolution predictors and high-resolution static data only (no noise addition is performed) and we, therefore, obtain a deterministic GAN.
@@
The discriminator is a PatchGAN convolutional classifier (Isola et al., 2017), which focuses on structures at the scale of image patches. The structure of the discriminator is composed of modules of the form convolution-BatchNorm-ReLu. It assigns a "realness" score to each N × N patch of the image, runs convolutionally across the image, and allows us to obtain an overall score by averaging all responses for each patch. High-quality results can be obtained with patches much smaller than the full size of the image, with relevant advantages in terms of resources for the training and application to arbitrarily large images. Details on the discriminator network structure and resolutions are depicted in Figure 3. The total number of trainable parameters for the GAN is $\sim$34 M (Table 1).
The training procedure follows the combined loss function approach for GANs (Goodfellow et al., 2020), including recent improvements to promote stability in the training (Esser et al., 2021), with the primary goal of balancing the minimization of both the generator’s and discriminator’s losses, which are adversarial. The pixel loss we used is the mean absolute error, while the discriminator loss is the hinge loss. The discriminator is activated after 50000 training steps, giving the generator time to learn to generate consistent outputs and thus stabilizing the adversarial training (Esser et al., 2021). After activating the discriminator, the network is trained by updating alternatively the gradients of the generator and the discriminator. The network is constantly fed with the whole target domain and no patch-training is applied.
@@ 3.3 Latent Diffusion Model Architecture
Diffusion Models (Sohl-Dickstein et al., 2015) are probabilistic models meant to extrapolate a data distribution p(x) by corrupting the training data through the successive addition of Gaussian noise (fixed) and then learning to recover the data by reversing this noising process (generative).
The Latent Diffusion Model (LDM) applied for this experiment is an architecture derived from Stable Diffusion (Rombach et al., 2021), specifically a re-adaptation of the conditional LDM LDCast (Leinonen et al., 2023), developed for precipitation nowcasting and already successfully applied for other variables (e.g., Carpentieri et al., 2023). The latent diffusion model derived for the downscaling task is composed of three main elements: (i) a convolutional Variational AutoEncoder (VAE), used to project the residual high-resolution target variables to and from a latent space; (ii) a conditioner, used to embed low-resolution data and high-resolution static data; and a denoiser, used to manage the diffusion process within the latent space. In the following sections, we report a detailed description of each model component. The total number of trainable parameters is shown in Table 1, while Figures 4 and 5 summarize the training and inference procedures for the model and the main structure of each component architecture, respectively.
@@ 3.3.1 Variational Autoencoder
The encoder and the decoder are structured as 2D convolutional networks composed of blocks of a ResNet residual block (Stephan et al., 2008) and a downsampling/upsampling convolutional layer. Three levels of such blocks are used, each reducing each spatial dimension by a factor of 2, while the number of channels is bottlenecked at 32 times the number of input target variables (i.e. 32x1 for the 2-m temperature and 32x2 for the 10-m wind speed components). The VAE bottleneck latent space is regularised with a loss based on Kullback-Leibler divergence (KL) (Kullback and Leibler, 1951; Csiszar, 1975) between the latent variable and a multivariate standard normal variable.
While the space dimensions are reduced by a factor of $8^{2}$ (from 512x512 to 64x64 pixel patches), the number of channels is increased from 1 (for the 2-m temperature) and 2 (for the 10-m wind speed components) to 32 and 64, respectively: the overall amount of data is therefore compressed only by a factor of 2, for both the VAEs (from 1x512x512 to 32x64x64 and from 2x512x512 to 64x64x64). Nevertheless, the gain in training performance of the denoiser and conditioner is much greater than the data reduction factor as the compression along the space dimension is more relevant for reducing the computational cost of the training than the increase in channel number (Rombach et al., 2021).
@@ 3.3.2 Conditioner
The conditioner stack acts as a context encoder to process the low-resolution predictors and high-resolution static data and embed them into each level of the denoiser UNET architecture. Initially, both datasets are preprocessed by passing through a dedicated encoder, a projection layer, and an analysis sequence, before being merged. The predictors’ encoder is a basic Identity layer as they already have the same spatial dimensions as the latent space (64x64). The static data encoder is a variational encoder with the same structure as the VAE described in the previous Section 3.3.1. For both datasets, the projection layer is a 2D convolutional layer with unitary kernel size, used to increase the number of channels, and the analysis is a sequence of 4 2D Adaptive Fourier Neural Operator (AFNO) blocks (following Pathak et al. (2022)), used to extract relevant features. After pre-processing, the conditioning information is prepared to be fed into each level of the denoiser UNET by applying a combination of average pooling and 2D Resnet layers. Figure 4 depicts details of the conditioner’s structure.
@@ 3.3.3 Denoiser
Our denoising stack is structured as the one of the LDCast (Leinonen et al., 2023), a re-adaptation of the U-Net-type network applied in the original latent diffusion model (Rombach et al., 2021). The resulting denoiser network consists of a UNET backbone enabled with a conditioning mechanism based on 2D AFNO blocks (Leinonen et al., 2023), aiming at a cross-attention-like operation (as suggested in Guibas et al. (2022)). This structure is meant to control the high-resolution synthesis process feeding the conditioning in each level of the UNET architecture.
@@
To improve the reconstruction of extreme values (for both temperature and wind speed) we implemented the v-prediction parameterization in our LDM model, following Salimans and Ho (2022): this parameterization trains the denoiser to model a weighted combination of both the noise and the start image, instead of either the only noise or the only start image as done in the more traditional implementations eps and x0, respectively.
As shown in Figure 5, the conditioner and the denoiser are trained together, minimizing the mean square error (L2), and feeding the network with random patches of ERA5 predictors (64x64 pixels) and static data (512x512 pixels) for the conditioning, and high-resolution target variables (512x512 pixels) for the ground truth. The training is performed using the AdamW optimizer (Loshchilov and Hutter, 2019), and Exponential Moving Averaging (EMA) is applied to the network weights, following Rombach et al. (2021).
@@ 4 Results
The following sections compare the results from LDM_res against the baselines using various verification metrics and distributions. In the Supporting Information we report the comparison of results from the LDM trained with and without the residual approach to provide an overview of the improvements introduced by this method.
@@ 4.1 Qualitative evaluation
To provide a qualitative and perceptual overview of the obtained results, we present a random snapshot of downscaled variables compared with both the input ERA5 low-resolution data and the COSMO_CLM high-resolution reference-truth (Figure 6). The second and third columns show a zoom-in on Sardinia Island, providing a deeper overview of models’ performance over complex terrain, coastal shores, and open sea. Both generative models, the GAN and LDM_res, effectively overcome the blurriness observed in both the quadratic interpolation and the UNET for the target variables. Particularly for 2-m temperatures, LDM_res demonstrates a remarkable ability to identify and reconstruct discontinuities in the variable field (zoomed-in view in Figure 6). Figure 6 also includes results for 10-meter wind speed (in color), which is a derived field obtained by combining the two actual target variables of the models, U and V. Perceptually, the results for this variable from both the GAN and LDM_res appear similar and equally plausible, displaying significantly more small-scale features compared to the UNET. A deeper qualitative examination reveals that the GAN aligns well with the reference truth, particularly over land, but exhibits mode collapse over the sea for both target variables. An example of this effect is shown in the Supporting Information. Conversely, the LDM_res consistently generates plausible high-resolution data across the entire domain, over both land and sea, and for both target variables.
Figure 6: Downscaled variables from all the tested models against low-resolution ERA5 input data and high-resolution COSMO_CLM reference truth, for a randomly picked timestamp. The left columns refer to 2-m temperature, and the right columns refer to 10-m wind speed. The second and fourth columns show a zoom-in on Sardinia Island.
@@ 4.2 Verification deterministic metrics
Figure 7 compares model results for different deterministic metrics, averaging results over the whole domain for each test timestep. In addition to results from the baseline and tested models, Figure 7 also reports results for the VAE of the LDM. These results are obtained using the VAE offline, feeding it with COSMO_CLM high-resolution data and calculating the metrics on the reconstructed data: this allows quantifying the LDM error resulting from the data decompression from the latent space only. We present three distance metrics, the Root Mean Square Error (RMSE), the mean bias (BIAS), and the coefficient of determination (R2), and one correlation metric, the Pearson Correlation Coefficient (PCC) (all calculated following Bell et al. (2021)).
@@
The results show that the VAE contribution to LDM_res is the highest for the RMSE of 2-m temperature, while bias, R2, and PCC have very little to no effects on temperature and wind speed.
@@ 4.3 Spatial distribution of errors
The spatial distribution of averaged-in-time magnitude differences for both the target variables and all tested models is illustrated in Figure 8. Within each panel, the numbers in squared brackets represent the 0.5 and 99.5 percentile values, offering insight into the highest errors recorded over the domain. Negative and positive values signify underestimation and overestimation, respectively, for both variables. Results from the quadratic interpolation of ERA5 data provide information on the original input data: 2-m temperature tends to be highly overestimated over complex terrain but underestimated on flat terrain, with smaller errors over the sea. Wind speed, conversely, is largely underestimated over land, particularly over mountain ridges, with a tendency towards overestimation along coastal shores.
@@
As for the wind speed results, all models exhibit a tendency to underestimation. LDM_res demonstrates superior performance, minimizing errors to nearly zero over most of the domain, with a uniform distribution over land and sea. The GAN displays traces of its characteristic mode collapses, especially over the sea: this evidence indicates that these mode collapses persist statically in fixed locations over time, consistently with the deployed training approach (i.e. feeding the network always the entire, fixed domain).
@@ 4.4 Frequency distributions
Figure 9: Comparison of frequency distributions for results from the tested models against COSMO_CLM reference-truth. The left column refers to the 2-m temperature, and the right column refers to the 10-m wind speed. Counting of pixel-wise data is cumulated for the yearly test dataset over bins of 0.5 °C and 0.05 m/s for temperature and wind speed, respectively. Notice that y-axes are logarithmic to highlight the tails of the distributions, hence the extreme values. The top row focuses on the tails of the distributions, i.e. on extreme values, while the bottom row focuses on the most frequent values and is a zoom-in on the dashed boxes for each variable.
@@
Reconstructing the distribution of 10-m wind speed proves more challenging for all models, given the inherent chaotic nature of the $U$ and $V$ wind components compared to temperature, which is strongly influenced by terrain elevation. Nonetheless, performance outcomes mirror those of the 2-m temperature. The GAN modestly improves upon UNET results, primarily in reducing occurrences of low wind speeds. LDM_res exhibits the highest performance in capturing both the tail and center of the wind speed distribution.
@@ 4.5 Radially Averaged Power Spectral Density (RAPSD)
Figure 10: Comparison of Radially Averaged Power Spectral Density (RAPSD) distributions for results from the tested models against COSMO_CLM reference-truth. The left column refers to the 2-m temperature, and the right column refers to the 10-m wind speed. The first row shows the averaged-in-time spectra, across the whole test dataset. Notice that in the first row y-axes are logarithmic to highlight the tail of the distributions, hence the high frequencies. The bottom rows show the distributions of single-time RAPSD values for fixed wavelengths, namely 269, 20, 9, and 5 km.
@@
Figure 10 showcases the results in terms of Radially Averaged Power Spectral Density (RAPSD), computed following the implementation outlined in Pulkkinen et al. (2019). The top row of the figure illustrates a single RAPSD, representing the average of each RAPSD calculated for every timestamp within the test dataset. To provide insight into the distribution of these values across all timestamps, the distributions of single-time RAPSD for fixed wavelengths are displayed in the bottom rows of Figure 10.
@@
Overall, all DL models effectively reconstruct the 2-m temperature power spectra down to wavelengths of 10 km. However, LDM_res consistently outperforms both the UNET and the GAN, as evident in panels (b) and (c) of Figure 10. The UNET and the GAN yield similar results, with marginal yet consistent enhancements originating from the adversarial training of the UNET. The diffusion process of LDM_res adeptly enhances the generation of small-scale features, showing precise reconstruction of the spectra up to 9 km (see panel c). For scales smaller than 9-10 km, all models exhibit decreased performance, albeit still showing improvements over the quadratic interpolation of ERA5. LDM_res still outperforms the other models but the very small-scale variability of the original data is slightly underestimated. This behavior is to be ascribed to the VAE, as further elucidated in the Supporting Information. Indeed, original reference-truth data compressed and reconstructed by the VAE show the very same power spectra underestimation for scales smaller than 9 km. The loss of information is therefore due to and inherent to the projection to the latent space.
@@
In contrast, results for wind speed distinctly demonstrate that generative models surpass both quadratic interpolation and UNET, effectively matching the slope of the energy power spectra and remaining competitive with each other. Specifically, LDM_res consistently outperforms the GAN up to 7 km, as emphasized in panels (b) and (c) of Figure 10 (right column). Similar to the 2-m temperature, for scales smaller than 10-9 km, both the GAN and LDM_res experience reduced performance, although they consistently exhibit improvements over the UNET. This behavior, for LDM_res, is in this case only partly to be ascribed to the VAE, as further shown in the Supporting Information, and an additional loss, for scales smaller than 9 km, is to be attributed intrinsically to the extraction of features with the diffusion process conditioned with the low-resolution data and high-resolution static data.